Evaluate the integral of sec 5θ tan 5θ dθ


1
Expert's answer
2021-03-10T11:20:01-0500

Let us evaluate the integral of "\int \sec 5θ \tan 5θ dθ:"


"\int \sec 5θ \tan 5θ dθ=\int \frac{1}{\cos 5θ}\frac{ \sin 5θ}{\cos 5θ} dθ=
-\frac{1}{5}\int \frac{1}{\cos^2 5θ} d(\cos 5θ)=
-\frac{1}{5}\int \cos^{-2} 5θ d(\cos 5θ)=


\frac{1}{5}\ \cos^{-1} 5θ + C=\frac{1}{5\cos 5θ}+C"